OpenStudy (anonymous):

Find the first two positive solutions. 7sin (π/5*X)=6

OpenStudy (anonymous):

sin(pi/5x)=6/7

OpenStudy (anonymous):

I found x= 5/pi Arcsin(6/7) . Is that right?

OpenStudy (anonymous):

then \[\frac{ \pi }{ 5x }=\sin^{-1} \frac{ 6 }{ 7 }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks